# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ I-D Contributions
Paper Organization: In Section II, we introduce the system model and formulate the problem. In Section III, we derive a CLT for the MID. Based on the CLT, we give the closed-form expressions for the upper and lower bounds of the average error probability in Section IV. In Section V, we provide the simulation results and Section VI concludes the paper.
Notations: Bold, upper case letters and bold, lower case letters represent matrices and vectors, respectively. $\mathbb{P}(\cdot)$ represents the probability operator and $\mathbb{E}x$ denotes the expectation of $x$. $\mathbb{C}^{N}$ and $\mathbb{C}^{M\times N}$ denote the $N$-dimensional vector space and the $M$-by-$N$ matrix space, respectively. $\mathbb{A}^{H}$ represents the conjugate transpose of $\mathbb{A}$, and the $(i,j)$-th entry of $\mathbb{A}$ is denoted by $[{\mathbb{A}}]_{i,j}$ or $A_{ij}$. $(\cdot)^{*}$ represents the conjugate of a complex number and $\|{\mathbb{A}}\|$ denotes the spectral norm of $\mathbb{A}$. $\operatorname{Tr}{\mathbb{A}}$ refers to the trace of ${\mathbb{A}}$, and $\mathbb{I}_{N}$ denotes the $N$ by $N$ identity matrix. The cumulative distribution function (CDF) of the standard normal distribution is denoted by $\Phi(x)$. $\underline{x}=x-\mathbb{E}x$ denotes the centered random variable and $\mathrm{Cov}(x,y)=\mathbb{E}\underline{x}\underline{y}$ represents the covariance between $x$ and $y$. $N\rightarrow\infty]{\mathcal{D}}$ denotes the convergence in distribution. $a\downarrow b$ represents that $a$ approaches $b$ from the right. $\mathcal{P}(x)$ denotes a polynomial with positive coefficients, and $\mathrm{supp}(\cdot)$ denotes the support operator. Given a set $\mathcal{S}$, $\mathrm{P}(\mathcal{S})$ denotes the set of probability measures with support of a subset of $\mathcal{S}$. The big-O, little-O, and big-theta notations are represented by ${\mathcal{O}}(\cdot)$ $o(1)$, and $\Theta(\cdot)$, respectively.
@@ III-A Preliminary Results
$C(\sigma^{2})=\frac{1}{M}\log\det(\mathbb{I}_{N}+\frac{1}{\sigma^{2}}{\mathbb{H}}{\mathbb{H}}^{H})$ (6)
@@
$\overline{C}(\sigma^{2})$ $=-\frac{\log(\sigma^{2})}{\kappa}-(\eta-\frac{1}{\kappa})\log(1-\frac{\omega}{\eta(1+\omega)})$ (9)
@@
$P(\sigma^{2})$ $=\omega^{3}+(2\sigma^{2}+\eta\kappa-\kappa-\eta+1)\frac{\omega^{2}}{\sigma^{2}}$ (10)
$+(1+\frac{\eta\kappa}{\sigma^{2}}-\frac{2\eta}{\sigma^{2}}+\frac{1}{\sigma^{2}})\omega-\frac{\eta}{\sigma^{2}}=0,$
@@
$\delta=\frac{1}{\sigma^{2}}(\frac{N}{L}-\frac{M\omega}{L(1+\omega)}),~{}~{}\overline{\omega}=\frac{1}{1+\omega}.$ (11)
@@ III-B CLT for the MID
×$\times(\frac{\omega^{2}(1+\delta\overline{\omega})}{1+\delta\overline{\omega}^{2}}-\frac{\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})}),$
@@
$\mathbb{P}(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x)=\Phi(x)+{\mathcal{O}}(n^{-\frac{1}{4}}).$ (19)
@@
$\omega^{2}+(1-\frac{\eta}{\sigma^{2}}+\frac{1}{\sigma^{2}})\omega-\frac{\eta}{\sigma^{2}}=0,$ (20)
@@
IV Applications to Bounds for the Optimal Average Error Probability
@@ IV-A Performance Metrics
The encoder mapping: A $(\mathrm{P}_{e}^{(n)},G)$-code for the system in (2) can be represented by the following mapping $f$,
@@
denotes the maximal energy constraint and $m$ is uniformly distributed in $\mathcal{C}_{n}=\{1,2,..,G\}$. Here $\mathcal{C}_{n}$ is the codebook, i.e., $\{f(1),f(2),...,f(G)\}$. There are two types of constraints on the channel inputs, i.e., the maximal energy constraint in (27) and the equal energy constraint (sphere constraint) in (14). Obviously, the inputs satisfying the equal energy constraint is a subset of those following the maximal energy constraint. Our goal is to obtain the bounds for the error probability with the maximal energy constraint in (27).
@@
$g:\mathbb{C}^{N\times n}\rightarrow\mathcal{G}\cup\{e\}.$ (28)
@@ A.1. Average Error Probability
$\mathrm{P}_{e}^{(n)}(\mathcal{C}_{n})=\frac{1}{G}\sum_{i=1}^{G}\mathbb{P}(\hat{m}\neq m|m=i),$ (29)
@@
$\mathrm{P}_{e}^{(n)}(R)=\inf_{\mathrm{supp}(\mathcal{C}_{n})\subseteq\mathcal{S}^{n}}\mathrm{P}_{e}^{(n)}(\mathcal{C}_{n}),$ (30)
@@ A.2. The Optimal Average Error Probability with Respect to the Second-Order Coding Rate
$\mathbb{P}_{e}(r|\rho,\eta,\kappa){=}\liminf\limits_{\mathrm{supp}(\mathcal{C}_{n})\subseteq\mathcal{S}^{n}}\limsup_{n\xrightarrow[]{\rho,\eta,\kappa}\infty}\mathrm{P}_{e}^{(n)}(\mathcal{C}_{n}),$ (32)
where $C(\sigma^{2})=\frac{1}{M}\log\det(\mathbb{I}_{N}+\frac{1}{\sigma^{2}}{\mathbb{H}}{\mathbb{H}}^{H})$ denotes the per antenna capacity and $\mathrm{P}_{e}^{(n)}(\mathcal{C}_{n})$ in (29) represents the average error probability of code $\mathcal{C}_{n}$. From (30) and (32), we can observe that for $r={\mathcal{O}}(1)$, the per antenna rate $R=\frac{\log(|\mathcal{C}_{n}|)}{Mn}=\mathbb{E}[C(\sigma^{2})]+\frac{r}{\sqrt{Mn}}$ is a ${\mathcal{O}}(\frac{1}{\sqrt{Mn}})$ perturbation around $\mathbb{E}[C(\sigma^{2})]$.
@@
$\mathbb{F}(r|\rho,\eta,\kappa)\leq\mathbb{P}_{e}(r|\rho,\eta,\kappa)\leq\mathbb{G}(r|\rho,\eta,\kappa),$ (33)
@@
Note that the $\inf$ operation is taken over $\{\mathbb{P}({{\mathbb{X}}^{(n+1)})}\in\mathrm{P}(\mathcal{S}_{=}^{n+1})\}$ in $\mathbb{F}(r|\rho,\eta,\kappa)$, where $\mathrm{P}(\mathcal{S}_{=}^{n+1})$ denotes the set of probability measures on codes with support a subset of $\mathcal{S}_{=}^{n+1}$. This results from the adaptation from the maximal energy constraint to the equal energy constraint by introducing an auxiliary symbol [7, Lemma 39]. Lemma 1 is important due to two reasons. First, it converts the evaluation with constraint $\mathcal{S}$ to that on the sphere coding (equal energy constraint) $\mathcal{S}_{=}$ in (14). Second, it indicates that, to characterize $\mathbb{P}_{e}(r|\rho,\eta,\kappa)$, we need to investigate the distribution of the MID. The parameter $\zeta$ will be used when analyzing the cases with $r>0$ and $r\leq 0$ in the proof of Theorem 3.
@@ B. FBL analysis
(Bounds for the optimal average error probability) The optimal average error probability $\mathbb{P}_{e}(r|\rho,\eta,\kappa)$ for the second-order coding rate over Rayleigh-product channels is bounded by
$\mathbb{P}_{e}(r|\rho,\eta,\kappa)\geq\begin{cases}\Phi(\frac{r}{\sqrt{V_{-}}})+{\mathcal{O}}(n^{-\frac{1}{2}}),~{}~{}r\leq 0,\\ \frac{1}{2},~{}~{}r>0,\end{cases}$ (35)
$\mathbb{P}_{e}(r|\rho,\eta,\kappa)\leq\Phi(\frac{r}{\sqrt{V_{+}}})+{\mathcal{O}}(n^{-\frac{1}{2}}),$ (36)
@@
$V_{+}$ $=V_{-}+\kappa\overline{\omega}^{4}[\frac{\omega^{2}(1+\delta\overline{\omega})}{1+\delta\overline{\omega}^{2}}-\frac{\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})}].$
@@
00 0 $<\Phi(\frac{r}{\sqrt{V_{+}}})-\Phi(\frac{r}{\sqrt{V_{-}}})\leq\frac{-r(V_{+}-V_{-})}{\sqrt{2\pi V_{+}V_{-}}(\sqrt{V_{+}}+\sqrt{V_{-}})}$ (38)
@@
$=\Phi(\frac{M(R-\overline{C}(\sigma^{2}))}{\sqrt{-\log(\Xi)}})\approx\mathrm{P}_{out}(\eta,\kappa,R),$
@@
$W(\omega)$ $=\log(\frac{(1+\omega)^{2}}{1+2\sigma^{2}\omega^{3}+2\sigma^{2}\omega^{2}}),~{}$ (42)
@@
$Y(\omega)$ $=\frac{\kappa}{(1+\omega)^{4}}[\frac{\omega^{2}(\sigma^{2}(1+\omega)^{2}+\omega+1+\kappa)}{\sigma^{2}(1+\omega)^{2}+(1-\kappa)\omega+1+\kappa}$
$+\frac{[1+(1-\kappa)\omega]^{4}X(\omega)}{\kappa^{2}[(1-\kappa)\omega^{2}+2\omega+1]^{2}}].$
@@ V-A Approximation Accuracy of Upper and Lower Bounds
Figure 4: Bounds with different numbers of the scatterers.
@@ V-C Impact of the Number of Scatterers on the Packet Error Probability
V-D Theoretical Bounds and Performance of LDPC
@@
where $L(s_{i}|\mathbb{r},{\mathbb{H}})$ represents the log likelihood ratio of the $i$-th bit $s_{i}\in\{0,1\}$ and $\mathcal{C}^{(i)}_{1}=\{\mathbb{c}|c_{i}=1,\mathbb{c}\in\mathcal{C}\}$ denotes the set of the codewords whose $i$-th digit is $1$. Here $\mathcal{C}^{(i)}_{0}$ represents the set of the codewords whose $i$-th digit is 00. $p(\mathbb{r}|\mathbb{c},{\mathbb{H}})$ is the conditional probability density function of the received signal $\mathbb{r}$. The output of the demodulator is then decoded by the soft-decision LDPC decoder. The length of the LDPC codes are $l\in\{567,2304\}$ bits, which correspond to the blocklengths $n=\frac{l}{2M}\in\{36,144\}$. The packet error is compared with the theoretical bounds in Theorem 3 for different SNRs ($\frac{1}{\sigma^{2}}$). The second-order coding rate is $r=\sqrt{Mn}(R-\overline{C}(\sigma^{2}))$. From Fig. 5, it can be observed that the upper bound and lower bound are nearly overlapped, which validates the tightness of the bounds. Similar to the single Rayleigh case [14], the theoretical bounds are nearly parallel with the error probability of the LDPC codes and the gap is around $2$ dB in the considered SNR range. Similar phenomenon can also be observed from the low rank case ($L=4$) but with worse performance, as shown in Fig. 6.

### 09 Apr 2024 v5

@@ I-D Contributions
Paper Organization: Section II introduces the MIMO system over Rayleigh-product channels and the definition of MID. Section III presents the CLT to characterize the asymptotic distribution of the MID. Section IV gives the the closed-form expressions for the upper and lower bounds of the average error probability. Section V shows the numerical results and Section VI concludes the paper.
Notations: Bold, upper case letters and bold, lower case letters represent matrices and vectors, respectively. The probability and expectation operator are denoted by $\mathbb{P}(\cdot)$ and $\mathbb{E}[\cdot]$, respectively. The $N$-dimensional vector space and $M$-by-$N$ matrix space are represented by $\mathbb{C}^{N}$ and $\mathbb{C}^{M\times N}$. The conjugate transpose, spectral norm, and trace of $\mathbb{A}$ are denoted by $\mathbb{A}^{H}$, $\|{\mathbb{A}}\|$, and $\operatorname{Tr}({\mathbb{A}})$, respectively. The $(i,j)$-th entry of $\mathbb{A}$ is denoted by $[{\mathbb{A}}]_{i,j}$ or $A_{ij}$. The conjugate of a complex number is represented by $(\cdot)^{*}$ and the $N$ by $N$ identity matrix is denoted by $\mathbb{I}_{N}$. The CDF of the standard normal distribution is denoted by $\Phi(x)$. The centered random variable is represented as $\underline{x}=x-\mathbb{E}[x]$ and the covariance between $x$ and $y$ is denoted by $\mathrm{Cov}(x,y)=\mathbb{E}[\underline{x}\underline{y}]$. The convergence in distribution is represented by $N\rightarrow\infty]{\mathcal{D}}$ and the limit as $a$ approaches $b$ from the right is denoted by $a\downarrow b$. The polynomial with positive coefficients and the support operator are represented by $\mathcal{P}(x)$ and $\mathrm{supp}(\cdot)$, respectively. Given a set $\mathcal{S}$, $\mathrm{P}(\mathcal{S})$ denotes the set of probability measures with support of a subset of $\mathcal{S}$. The big-O, little-O, and big-theta notations are represented by ${\mathcal{O}}(\cdot)$, $o(1)$, and $\Theta(\cdot)$, respectively.
@@ III-A Preliminary Results
$C(\sigma^{2})=\frac{1}{M}\log\det\left(\mathbb{I}_{N}+\frac{1}{\sigma^{2}}{\mathbb{H}}{\mathbb{H}}^{H}\right)$ (6)
@@
$\overline{C}(\sigma^{2})$ $=-\frac{\log(\sigma^{2})}{\kappa}-\left(\eta-\frac{1}{\kappa}\right)\log\left(1-\frac{\omega}{\eta(1+\omega)}\right)$ (9)
@@
$P(\sigma^{2})$ $=\omega^{3}+\left(2\sigma^{2}+\eta\kappa-\kappa-\eta+1\right)\frac{\omega^{2}}{\sigma^{2}}$ (10)
$+\left(1+\frac{\eta\kappa}{\sigma^{2}}-\frac{2\eta}{\sigma^{2}}+\frac{1}{\sigma^{2}}\right)\omega-\frac{\eta}{\sigma^{2}}=0,$
@@
$\delta=\frac{1}{\sigma^{2}}\left(\frac{N}{L}-\frac{M\omega}{L(1+\omega)}\right),~{}~{}\overline{\omega}=\frac{1}{1+\omega}.$ (11)
@@ III-B CLT for the MID
×$\times\left(\frac{\omega^{2}(1+\delta\overline{\omega})}{1+\delta\overline{\omega}^{2}}-\frac{\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})}\right),$
@@
$\mathbb{P}\left(\!\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\!\leq\!x\!\right)\!=\!\Phi(x)+{\mathcal{O}}(n^{-\frac{1}{4}}).$ (19)
@@
$\omega^{2}+\left(1-\frac{\eta}{\sigma^{2}}+\frac{1}{\sigma^{2}}\right)\omega-\frac{\eta}{\sigma^{2}}=0,$ (20)
@@
IV Bounds for the Optimal Average Error Probability
@@ IV-A Performance Metrics
The encoder mapping: A $(\mathrm{P}_{\mathrm{e}}^{(n)},G)$-code for the system in (2) can be represented by the following mapping $f$,
@@
denotes the maximal energy constraint and $m$ is uniformly distributed in $\mathcal{G}=\{1,2,..,G\}$. Here $\mathcal{C}_{n}$ is the codebook, i.e., $\{f(1),f(2),...,f(G)\}$. There are two types of constraints on the channel inputs, i.e., the maximal energy constraint in (27) and the equal energy constraint (sphere constraint) in (14). Obviously, the inputs satisfying the equal energy constraint is a subset of those following the maximal energy constraint. Our goal is to obtain the bounds for the error probability with the maximal energy constraint in (27).
@@
$g:\mathbb{C}^{N\times n}\rightarrow\mathcal{G}\cup\{\mathrm{e}\}.$ (28)
@@ A.1. Average Error Probability
$\mathrm{P}_{\mathrm{e}}^{(n)}(\mathcal{C}_{n})=\frac{1}{G}\sum_{i=1}^{G}\mathbb{P}(\hat{m}\neq m|m=i),$ (29)
@@
$\mathrm{P}_{\mathrm{e}}^{(n)}(R)=\inf_{\mathrm{supp}(\mathcal{C}_{n})\subseteq\mathcal{S}^{n}}\mathrm{P}_{\mathrm{e}}^{(n)}(\mathcal{C}_{n}),$ (30)
@@ A.2. The Optimal Average Error Probability with Respect to the Second-Order Coding Rate
$\mathbb{P}_{\mathrm{e}}(r|\rho,\eta,\kappa){=}\liminf\limits_{\mathrm{supp}(\mathcal{C}_{n})\subseteq\mathcal{S}^{n}}\limsup_{n\xrightarrow[]{\rho,\eta,\kappa}\infty}\mathrm{P}_{\mathrm{e}}^{(n)}(\mathcal{C}_{n}),$ (32)
where $C(\sigma^{2})=\frac{1}{M}\log\det(\mathbb{I}_{N}+\frac{1}{\sigma^{2}}{\mathbb{H}}{\mathbb{H}}^{H})$ denotes the per antenna capacity and $\mathrm{P}_{\mathrm{e}}^{(n)}(\mathcal{C}_{n})$ in (29) represents the average error probability of code $\mathcal{C}_{n}$. From (30) and (32), we can observe that for $r={\mathcal{O}}(1)$, the per antenna rate $R=\frac{\log(|\mathcal{C}_{n}|)}{Mn}=\mathbb{E}[C(\sigma^{2})]+\frac{r}{\sqrt{Mn}}$ is a ${\mathcal{O}}(\frac{1}{\sqrt{Mn}})$ perturbation around $\mathbb{E}[C(\sigma^{2})]$.
@@
$\mathbb{F}(r|\rho,\eta,\kappa)\leq\mathbb{P}_{\mathrm{e}}(r|\rho,\eta,\kappa)\leq\mathbb{G}(r|\rho,\eta,\kappa),$ (33)
@@
Note that the $\inf$ operation is taken over $\{\mathbb{P}({{\mathbb{X}}^{(n+1)})}\in\mathrm{P}(\mathcal{S}_{=}^{n+1})\}$ in $\mathbb{F}(r|\rho,\eta,\kappa)$, where $\mathrm{P}(\mathcal{S}_{=}^{n+1})$ denotes the set of probability measures on codes with support a subset of $\mathcal{S}_{=}^{n+1}$. This results from the adaptation from the maximal energy constraint to the equal energy constraint by introducing an auxiliary symbol [7, Lemma 39]. Lemma 1 is important due to two reasons. First, it converts the evaluation with constraint $\mathcal{S}$ to that on the sphere coding (equal energy constraint) $\mathcal{S}_{=}$ in (14). Second, it indicates that, to characterize $\mathbb{P}_{\mathrm{e}}(r|\rho,\eta,\kappa)$, we need to investigate the distribution of the MID. The parameter $\zeta$ will be used when analyzing the cases with $r>0$ and $r\leq 0$ in the proof of Theorem 3.
@@ B. FBL analysis
(Bounds for the optimal average error probability) The optimal average error probability $\mathbb{P}_{\mathrm{e}}(r|\rho,\eta,\kappa)$ for the second-order coding rate over Rayleigh-product channels is bounded by
$\mathbb{P}_{\mathrm{e}}(r|\rho,\eta,\kappa)\geq\begin{cases}\Phi(\frac{r}{\sqrt{V_{-}}})+{\mathcal{O}}(n^{-\frac{1}{2}}),~{}~{}r\leq 0,\\ \frac{1}{2},~{}~{}r>0,\end{cases}$ (35)
$\mathbb{P}_{\mathrm{e}}(r|\rho,\eta,\kappa)\leq\Phi(\frac{r}{\sqrt{V_{+}}})+{\mathcal{O}}(n^{-\frac{1}{2}}),$ (36)
@@
$V_{+}$ $=V_{-}+\kappa\overline{\omega}^{4}\left[\frac{\omega^{2}(1+\delta\overline{\omega})}{1+\delta\overline{\omega}^{2}}-\frac{\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})}\right].$
@@
00 0 $<\Phi\Bigl{(}\frac{r}{\sqrt{V_{+}}}\Bigr{)}-\Phi\Bigl{(}\frac{r}{\sqrt{V_{-}}}\Bigr{)}\leq\frac{-r(V_{+}-V_{-})}{\sqrt{2\pi V_{+}V_{-}}(\sqrt{V_{+}}+\sqrt{V_{-}})}$ (38)
@@
$=\Phi\Biggl{(}\frac{M(R-\overline{C}(\sigma^{2}))}{\sqrt{-\log(\Xi)}}\Biggr{)}\approx\mathrm{P}_{\mathrm{out}}(\eta,\kappa,R),$
@@
$W(\omega)$ $=\log\left(\frac{(1+\omega)^{2}}{1+2\sigma^{2}\omega^{3}+2\sigma^{2}\omega^{2}}\right),~{}$ (42)
@@
$Y(\omega)$ $=\frac{\kappa}{(1+\omega)^{4}}\Biggl{[}\frac{\omega^{2}(\sigma^{2}(1+\omega)^{2}+\omega+1+\kappa)}{\sigma^{2}(1+\omega)^{2}+(1-\kappa)\omega+1+\kappa}$
$+\frac{[1+(1-\kappa)\omega]^{4}X(\omega)}{\kappa^{2}[(1-\kappa)\omega^{2}+2\omega+1]^{2}}\Biggr{]}.$
@@ V-B Impact of $n$ on the Optimal Average Error Probability
Figure 4: Bounds with different numbers of the scatterers.
@@ V-C Impact of the Number of Scatterers on the Packet Error Probability
V-D Comparison Between Theoretical Bounds and Performance of LDPC
where $L(s_{i}|\mathbb{r},{\mathbb{H}})$ represents the log likelihood ratio of the $i$-th bit $s_{i}\in\{0,1\}$ and $\mathcal{C}^{(i)}_{1}=\{\mathbb{c}|c_{i}=1,\mathbb{c}\in\mathcal{C}\}$ denotes the set of the codewords whose $i$-th digit is $1$. Here $\mathcal{C}^{(i)}_{0}$ represents the set of the codewords whose $i$-th digit is 00. $p(\mathbb{r}|\mathbb{c},{\mathbb{H}})$ is the conditional probability density function of the received signal $\mathbb{r}$. The output of the demodulator is then decoded by the soft-decision LDPC decoder. The length of the LDPC codes are $l\in\{576,2304\}$ bits, which correspond to the blocklengths $n=\frac{l}{2M}\in\{36,144\}$. The packet error is compared with the theoretical bounds in Theorem 3 for different SNRs ($\frac{1}{\sigma^{2}}$). The second-order coding rate is $r=\sqrt{Mn}(R-\overline{C}(\sigma^{2}))$. From Fig. 5, it can be observed that the upper bound and lower bound are nearly overlapped, which validates the tightness of the bounds. Similar to the single Rayleigh case [14], the theoretical bounds are nearly parallel with the error probability of the LDPC codes and the gap is around $2$ dB in the considered SNR range. Similar phenomenon can also be observed from the low rank case ($L=4$) but with worse performance, as shown in Fig. 6.